An aqueous solution is prepared to be 0.381 M in sodium acetylsalicylate and 0.250 M in benzoic acid. (1) Is this solution a buffer solution? (2) What is the pH of this solution? pH = (3) If 0.145 moles of hydroiodic acid are added to one liter of this solution, what is the pH of the resulting solution? pH = Use the Tables link in the References for any equilibrium constants that are required.
Answer #1

1) A buffer solution is made up of an acid and its conjugate base, in this case we are NOT in the presence of a buffer, since the salt is not the conjugate pair of the acid.

2) The hydrolysis reaction of the acid is:

HA + H2O = A- + H3O +

From the expression of Ka we have:

Ka = [A-] * [H3O +] / [HA]

We replace:

6.46x10 ^ -5 = X ^ 2 / 0.250 - X

We cleared:

X ^ 2 + 6.46x10 ^ -5 X - 1.62x10 ^ -5 = 0

We apply equation of the second degree and:

X = [H3O +] = 4x10 ^ -3 M

Since the dissociation of the acetylsalicylate ion is very low, it is disregarded and pH is calculated:

pH = - Log [H3O +] = - Log (4x10 ^ -3) = 2.40

3) If 0.145 moles of perchloric acid are added to one liter of solution, it is assumed that a concentration of hydronium ion of 0.145 M is added, the new concentration of hydronium and the pH are calculated:

[H3O +] = 4x10 ^ -3 + 0.145 = 0.149 M

pH = - Log (0.106) = 0.827

If you liked the answer, please rate it in a positive way, you would help me a lot, thank you.